A particle moves on the path of 20 ft radius so that its arc distance from a fixed point on the path is given by s= 4³-10t where s is in feet and t in seconds. Compute the total acceleration at the end of 2 sec.

1
Expert's answer
2022-01-10T17:03:01-0500

s = 4t3 - 10t

v = ds/dt = 12t2 - 10

at​ = dv/dt = 24t

(at)t=2​ = 48

aR​ = V2/R = (Vt=2)2/20 = 482/20 = 115.2

​​a = (at2 + aR2)1/2 = (482 + 115.22) = 124.8 ft/sec